OpenStudy (anonymous):

What is the number of real solutions? 12x^2-11x=11

OpenStudy (jack1):

hi helen, sorri i was just finishing on another q

OpenStudy (jack1):

12x^2 - 11x = 11 so rearrange to get the quadratic equation y = ax^2 + bx + c 0 = 12x^2 - 11x - 11 u with me so far?